OpenStudy (anonymous):

Find the slope-intercept form for the line satisfying the following conditions. x-intercept=2, y intercept =4/5

OpenStudy (pradipgr817):

write it in the form (x/a)+(y/b)=1, then convert it into the form y=mx+c, you have your slope

OpenStudy (pradipgr817):

m is the slope

OpenStudy (pradipgr817):

a and b are intercept on x and y axis respectively.

OpenStudy (anonymous):

i still do not get it can you explain?

OpenStudy (pradipgr817):

you can write a equation in intercept form, if a is intercept on x axis and b on y, than you can write a line as (x/a)+(y/b)=1 is this point clear?

OpenStudy (anonymous):

nope this is all confusing to me i am horrible with math

jimthompson5910 (jim_thompson5910):

Slope intercept form is y = mx+b where b is the y-intercept

jimthompson5910 (jim_thompson5910):

Since the y intercept is given to be 4/5, we can say b = 4/5 and y = mx+b becomes y = mx + 4/5

jimthompson5910 (jim_thompson5910):

The x-intercept is given to be x = 2, so essentially, x = 2 and y = 0. Plug these values in and solve for m y = mx + 4/5 0 = m(2) + 4/5 0 = 2m + 4/5 0 - 4/5 = 2m + 4/5 - 4/5 -4/5 = 2m 2m = -4/5 m = (-4/5)*(1/2) m = -4/10 m = -2/5 This means that the equation of the line in slope intercept form is \(\Large y = -\frac{2}{5}x + \frac{4}{5}\)

OpenStudy (anonymous):

so that is the final answer there is no more simplifying or anything

jimthompson5910 (jim_thompson5910):

That's as simplified as it gets in that form

OpenStudy (anonymous):

did you divide by two towards the end, im just trying to make sure that i understand the process

jimthompson5910 (jim_thompson5910):

i multiplied both sides by 1/2, which is the same as dividing both sides by 2, so yes

jimthompson5910 (jim_thompson5910):

this was done to completely isolate m at the end

OpenStudy (anonymous):

okay so if i messed up on typing in the problem how much would this affect the answer? I should just be able to plug the right number in right ?

jimthompson5910 (jim_thompson5910):

You should be able to plug in each intercept and test if it works or not ex: plug in the x-intercept: x = 2 and y = 0 and you should get a true equation also plug in the y-intercept: plug in x = 0 and y = 4/5 and you should get a true equation

OpenStudy (anonymous):

okay i plugged in three and then i have y=-4/15x+4/5

jimthompson5910 (jim_thompson5910):

you multiply 3 by the numerator, not the denominator

jimthompson5910 (jim_thompson5910):

and once you plug in 3 for x, the x goes away

OpenStudy (anonymous):

okay im confused again lol i thought that was slope intercept form

jimthompson5910 (jim_thompson5910):

\(\Large y = -\frac{2}{5}x + \frac{4}{5}\) is slope intercept form

jimthompson5910 (jim_thompson5910):

and you're done because it cannot be simplified any more

OpenStudy (anonymous):

yeah okay but that was for the problem i messed up on and typed the wrong number into. so now im struggling with the problem once i am supposed to multiple by 1/3

OpenStudy (anonymous):

that is where i got the -4/15

jimthompson5910 (jim_thompson5910):

oh so the problem at the very top is incorrect?

OpenStudy (anonymous):

yeah i typed the wrong x intercept it was 3 not 2

jimthompson5910 (jim_thompson5910):

if the x-intercept is 3 instead of 2, then you are correct in saying that the answer is \(\Large y = -\frac{4}{15}x + \frac{4}{5}\)

OpenStudy (anonymous):

okay i was thrown off a little because of the denominators cuz 15 is divisible by 5

jimthompson5910 (jim_thompson5910):

i see

jimthompson5910 (jim_thompson5910):

well you definitely got the right answer because \(\Large y = -\frac{4}{15}x + \frac{4}{5}\) is fully simplified and it does have an x-intercept of 3 and a y-intercept of 4/5

OpenStudy (anonymous):

if the y intercept was 2/5s adn the x intercept was three would the solution be -4/15x+2/5